OpenStudy (anonymous):

Simplify this given expression to rational exponent form, justify each step by identifying the properties of rational exponents used.

OpenStudy (anonymous):



OpenStudy (anonymous):

My Answer: 1/x^(3/(-6)) I turned the radical in the denominator into a rational exponent. x^(3/6)/1 – I took the reciprocal of the above fraction. x^(1/2) – I simplified

OpenStudy (anonymous):

Is this right?

jimthompson5910 (jim_thompson5910):

it should be -6/3 = -2 in the exponent and NOT -3/6 = -1/2

jimthompson5910 (jim_thompson5910):

so we then will have \[\large \frac{1}{x^{-2}}\] which becomes \[\large x^2\]

OpenStudy (anonymous):

Wouldn't I just flip the whole equation around and the exponent stay the same?

jimthompson5910 (jim_thompson5910):

no the exponent will become positive once you flip the fraction

OpenStudy (anonymous):

Okay, thanks, I see what I did wrong now.

jimthompson5910 (jim_thompson5910):

you're welcome